MOMENTUM

What's the difference between m1v2 = m1v2 & m1v1 + m2v2 = m1v1' + m2v2' ?

When do you know which one to use and can you provide an example?
Answer #1

Concept of momentum and velocity is explained below :

let two separate masses are moving in a system.

let mass of a particle moving with velocity v1 is = m1

let mass of a particle moving with velocity v2 is= m2

Linear momentum = p = mass * velocity

as we know that when no resultant force acts on system of interacting particles the total momentum of system is conserved,therefore:

m1v1=m2v2

If both masses are moving with initial velocity v1, v2 and collision occurs between both particles then after collision let we assume that both particles are moving with speed v1', v2'

Then by principal of momentum conservation

m1v1+m2v2=m1v1'+m2v2' (total momentum before collision = total momentum after collision)

example:

a particle of mass 10kg moving with speed 5m/s collides with a second particle of mass 6kg travelling with speed 3m/s .If the speed of first particle after collision is 1.5m/s,then find the speed of second after collision

so in these situation you need to use the second relation of momentum conservation

m1v1+m2v2=m1v1'+m2v2' (total momentum before collision = total momentum after collision)

implies, 10*5 +6*3 =6*1.5 +6*v2

v2=9.833 m/s